OpenStudy (anonymous):

What is the area of a triangle with a base of 7 units and a height of 11 units? If necessary, round your answer to two decimal places.

OpenStudy (anonymous):

38.5 units

OpenStudy (anonymous):

A = 1/2 x Base x Height = 1/2 x 11 x 7

OpenStudy (anonymous):

Wht? got it?

OpenStudy (anonymous):

yeah i got the area now..